OpenStudy (anonymous):

dentify the correct slope and y-intercept of the equation 5x - 3y = -6

OpenStudy (anonymous):

y=mx+b -3y=-6-5x y=2-(5/3)x slope =-5/3 y intercept = 2

OpenStudy (anonymous):

thank you(: